OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (10, -4) and has slope -3/10. Please show all of your work

OpenStudy (anonymous):

general form of the equation is y=mx + b here m = -3/10, x = 10 and y = -4 substituting them in general form we get 3 -4 = - --- * 10 + b 10 -4 = -3 + b -4 + 3 = b -1 = b so the required equation is 3 y = - ---- x -1 10